# Restricted Liouville Operator for the study of Non-Analytic Dynamics within the Disk

### v1

@@ Abstract
Non-linear dynamics, System Identification, Liouville Operators, Occupation kernel, Hardy Spaces over Unit Disk.
Sushant Pokhriyal*, Joel A. Rosenfeld Corresponding author: sushantpokhriyal@gmail.com.
Department of Mathematics and Statistics, University of South Florida,
4202 E Fowler Avenue, Tampa, USA
sushantpokhriyal@gmail.com, rosenfeldj@usf.edu
@@ 1  Introduction
The study of nonlinear dynamical systems through linear operators can be traced back to Koopman and von Neumann’s work in the 1930s [11, 12]. More recently, the introduction of Dynamic Mode Decompositions (DMD) by Schmidt and Sesterhenn in 2008, [20], opened the door to data-driven analysis of nonlinear dynamical systems through linear operators. Their work was later expanded upon through extended DMD (EDMD) by Williams et al in [23], as a method that combines the selection of nonlinear features with DMD was extended to infinite feature spaces, and since then, reproducing kernel Hilbert spaces (RKHSs) have been an integral part of the study of DMD . Mezic and Rowley later connected the study of DMD with ergodic theory and the Koopman operator in [2, 15, 24]. For detailed review of these techniques, please refer to [1, 5], and references therein.
@@
In Gonzalez et al. [9], it was observed that by combining boundedness assumptions on Koopman operators together with regularity conditions of members of an RKHS, very strong conditions on the dynamics are imposed. For instance, all bounded Koopman operators over the Bargmann-Fock space have been shown to have affine symbols (dynamics) in Carswell et al. [3]. This same degeneracy also occurs when working over the native space of a Gaussian radial basis function, the RKHS of polynomials, etc. This motivates the use of other operators to represent non-linear dynamics over RKHSs, such as the Koopman generator (or more generally, Liouville operators). While Koopman generators presuppose forward completeness of nonlinear dynamics so that discrete-time dynamics may be generated, Liouville operators do not have this requirement but have the same formal definition as Koopman generators [17].
@@
While often unbounded, Liouville operators are densely defined for large classes of dynamics [18]. In Rosenfeld et al. [17], scaled Liouville operators were introduced, which are compact for a large collection of dynamics, and in Rosenfeld et al., the singular DMD method was introduced to provide a collection of compact Liouville operators without a scaling factor. The compactness guarantees are important in these settings, since the DMD method approximates dynamic operators through finite rank approximations derived from the available trajectory data of the non-linear dynamical systems.
@@
Recently, work by Mezic and Mauroy have been focused on the development of Koopman operators corresponding to analytic dynamics over the disc through the Hardy space [14], and work by Russo and Rosenfeld has begun to study Liouville operators over the Hardy space [19]. Despite the expansion of the dynamics considered admissible for DMD methods through Liouville operators, the dynamics are required to match the regularity properties of the functions of the spaces that they operate over [cite]. The removal of this limitation will allow for the consideration of a wider class of dynamics, while also retaining many of the benefits of RKHSs, such as bounded point evaluations. In the case of dynamical system which involves some non-analyticity, we will carefully construct a Hilbert space over which the Liouville operator is well defined. Specifically, we will define a symbol in $H^{\infty}$, that nullify the effect of non-analytic part of the dynamical system.
@@ 2  Motivation
Fortunately, we are not left completely in the dark. There are ways to make a non-linear dynamical system into linear systems, at the cost of moving from the finite dimensional setting to infinite dimensional one. In general, this can be achieved by looking at the dynamics under the action of some suitable linear operator (possibly unbounded), like Koopman, Perron-Frobeinous, Liouville, Carleman Linearization, etc, on some Hilbert space of functions over the state space $X$ [cite]. In all of these methods, we seek to approximate the unknown dynamics $f$ through the projection of $f$ over a set of suitable basis functions, which are mostly analytic in the Hilbert space on $X$. This forces regularity condition on $f$.
Consider the following example of two-compartment model to describe the drug distribution (Example 7 in [21]);
@@
There are important class of non-linear systems, where $f$ is beyond polynomials or analytic functions, for example, see Example 6 and 7 of drug distribution in [21], enzyme kinematics [10] or biochemical reactors [22], and references therein. In general, identification of such systems can be challenging [6]. In addition to this, problems in control can also give rise to system of differential equations in which the function $f$ can have discontinuities like finite jumps (switching on or off at certain time) or sudden changes in the behavior of the system [4]. Once during a test or simulation if a point of discontinuity is located, then our goal is to design the learning algorithm in such a way that it takes care of such discontinuous points by associating the symbol $f$ with suitable Hilbert space (See Figure 1). It is our desire to explore this avenue and start our quest of deep analysis towards this direction. Although, it is always not possible to figure out all such points during simulations, nonetheless we will start our investigation assuming that such points can be traced. The following items are the main contribution of this paper;
@@ 3.1  Reproducing Kernel Hilbert Space: RKHS
It is the closed subspace of $L^{2}(\mathbb{T})$. Vice versa, we can define the functions in the Hardy-Hilbert space $H^{2}(\mathbb{D})$, as the integral of the functions in $H^{2}(\mathbb{T})$ against the Poisson Kernel [7]. The reproducing kernel of $H^{2}(\mathbb{D})$ is known as the Szego Kernel and it is given by
@@
Kernel functions interact nicely with many classical function theoretic operators. For example, if $f:X\to\mathbb{C}$ is the symbol for a densely defined multiplication operator, $M_{f}:\mathcal{D}(M_{f})\to H$, with $\mathcal{D}(M_{f}):=\{g\in H:fg\in H\}$ and $M_{f}g:=gf$, then $M_{f}^{*}k_{x}=\overline{f(x)}k_{x}$. That is kernel functions are eigenfunctions for the adjoints of densely defined multiplication operators [16]. Similarly, the adjoint of a densely defined Koopman or composition operator, $\mathcal{K}_{\phi}$, on a kernel function is described as $\mathcal{K}_{\phi}^{*}k_{x}=k_{\phi(x)}$.
@@ 3.2  Liouville Operator and Occupation Kernel
It is easy to check that $A_{f}$ is a linear operator. Being a differential operator, $A_{f}$ is expected to be unbounded operator over $H^{2}(\mathbb{D})$, but as differential operators over RKHS’s consisting of continuously differentiable functions, it is always closed [18], hence $A_{f}$ is a closed operator over $H^{2}(\mathbb{D)}$.
@@
The above relation can be used to find finite approximation of the Liouville operator. For more information, we refer to the papers [18].
@@ 4  Restricted Liouville Operator
In this paper, we will pose the above problem inside $H^{2}(\mathbb{D})$, by taking $f$ as a symbol which is not analytic or smooth in nature. The following theorem from [19], explains the behavior of the symbol $f$ of the operator $A_{f}$, when $\overline{D(A_{f})}=H^{2}(\mathbb{D})$.
@@
Then it can be easily proved that $B_{j}\in H^{\infty}(\mathbb{D})$ and $|B_{j}(e^{i\theta})|=1$ for all $e^{i\theta}\in\mathbb{T}$. In literature, the function $B_{j}$’s are known as the finite Blaschke factors. One can easily show that $\|B_{j}g\|_{H^{2}}=\|g\|_{H^{2}}$ for all $g\in H^{2}$, and hence multiplication operator $T_{B_{j}}:H^{2}\rightarrow H^{2}$ is an isometry, which implies that the range space $T_{B_{i}}H^{2}=B_{i}H^{2}$ is an closed subspace of $H^{2}$ (see, [13]).
@@
Moreover, it is easy to see that $B\in H^{\infty}(\mathbb{D})$ and $|B(e^{i\theta})|=1$ for all $e^{i\theta}\in\mathbb{T}$ (see, [7]). Now, we are in the position to construct an Hilbert space which captures the singularities of the function $f$. Define the set
@@
Then $B^{2}H^{2}$ is a closed subspace of $H^{2}$ with $\{B^{2}z^{n}\}_{n=0}^{\infty}$ as its orthonormal basis [13]. Moreover, according to [8], the space $B^{2}H^{2}$ is an RKHS with the reproducing kernel defined as
@@ 6  Spectrum of Restricted Liouville Operator
Note, the symbol $f(z)=1/z$, has no zeroes inside the disk and has a pole inside the disk. On the other hand, according to Proposition 5 in [19], if $f$ is an analytic function with no zeroes in a neighborhood of the closed disk, then every $\lambda\in\mathbb{C}$ is in the point spectrum with $g(z)=C~\text{exp}~\Big(\int_{0}^{z}\frac{\lambda}{f(w)}dw\Big)$ as the corresponding eigenfunction. We can further improve Proposition 6.1, in the following way.
@@
    Spectrum of $A_{1/z}$ is continuous. Combining Proposition 6.1 and Theorem 6.2, we can conclude that the spectrum of the operator $A_{1/z}$ is whole complex plane, i.e., $\sigma(A_{1/z})=\mathbb{C}$. This is in contrast with the case $A_{z}$ where $\sigma(A_{z})=\mathbb{C}$, where the spectrum consists of only eigenvalues [Proposition 5, [19]]. Hence, the spectrum of the operator $A_{f_{|B^{2}H^{2}}}$ needs further investigations.

### v2

@@ Abstract
Non-linear dynamics , System Identification , Liouville Operators , Occupation kernel , Hardy Spaces over Unit Disk
@@ 1 Introduction
The study of nonlinear dynamical systems through linear operators can be traced back to Koopman and von Neumann’s work in the 1930s [koopman1931hamiltonian, koopman1932dynamical]. More recently, the introduction of Dynamic Mode Decompositions (DMD) by Schmidt and Sesterhenn in 2008, [schmid2010dynamic], opened the door to data-driven analysis of nonlinear dynamical systems through linear operators. Their work was later expanded upon through extended DMD (EDMD) by Williams et al in [williams2015data], as a method that combines the selection of nonlinear features with DMD was extended to infinite feature spaces, and since then, reproducing kernel Hilbert spaces (RKHSs) have been an integral part of the study of DMD . Mezic and Rowley later connected the study of DMD with ergodic theory and the Koopman operator in [budivsic2012applied, mezic2020spectrum, williams2014kernel]. For detailed review of these techniques, please refer to [brunton2022modern, colbrook2024multiverse], and references therein.
@@
In Gonzalez et al. [gonzalez2024kernel], it was observed that by combining boundedness assumptions on Koopman operators together with regularity conditions of members of an RKHS, very strong conditions on the dynamics are imposed. For instance, all bounded Koopman operators over the Bargmann-Fock space have been shown to have affine symbols (dynamics) in Carswell et al. [carswell2003composition]. This same degeneracy also occurs when working over the native space of a Gaussian radial basis function, the RKHS of polynomials, etc. This motivates the use of other operators to represent non-linear dynamics over RKHSs, such as the Koopman generator (or more generally, Liouville operators). While Koopman generators presuppose forward completeness of nonlinear dynamics so that discrete-time dynamics may be generated, Liouville operators do not have this requirement but have the same formal definition as Koopman generators [rosenfeld2023singular].
@@
While often unbounded, Liouville operators are densely defined for large classes of dynamics [rosenfeld2024occupation]. In Rosenfeld et al. [rosenfeld2023singular], scaled Liouville operators were introduced, which are compact for a large collection of dynamics, and in Rosenfeld et al., the singular DMD method was introduced to provide a collection of compact Liouville operators without a scaling factor. The compactness guarantees are important in these settings, since the DMD method approximates dynamic operators through finite rank approximations derived from the available trajectory data of the non-linear dynamical systems.
@@
Recently, work by Mezic and Mauroy have been focused on the development of Koopman operators corresponding to analytic dynamics over the disc through the Hardy space [mauroy2024analytic], and work by Russo and Rosenfeld has begun to study Liouville operators over the Hardy space [russo2022liouville]. Despite the expansion of the dynamics considered admissible for DMD methods through Liouville operators, the dynamics are required to match the regularity properties of the functions of the spaces that they operate over [cite]. The removal of this limitation will allow for the consideration of a wider class of dynamics, while also retaining many of the benefits of RKHSs, such as bounded point evaluations. In the case of dynamical system which involves some non-analyticity, we will carefully construct a Hilbert space over which the Liouville operator is well defined. Specifically, we will define a symbol in $H^{\infty}$, that nullify the effect of non-analytic part of the dynamical system.
@@ 2 Motivation
Fortunately, we are not left completely in the dark. There are ways to make a non-linear dynamical system into linear systems, at the cost of moving from the finite dimensional setting to infinite dimensional one. In general, this can be achieved by looking at the dynamics under the action of some suitable linear operator (possibly unbounded), like Koopman, Perron-Frobeinous, Liouville, Carleman Linearization, etc, on some Hilbert space of functions over the state space $X$. In all of these methods, we seek to approximate the unknown dynamics $f$ through the projection of $f$ over a set of suitable basis functions, which are mostly analytic in the Hilbert space on $X$. This forces regularity condition on $f$.
Consider the following example of two-compartment model to describe the drug distribution (Example 7 in [strasser2021data]);
@@
There are important class of non-linear systems, where $f$ is beyond polynomials or analytic functions, for example, see Example 6 and 7 of drug distribution in [strasser2021data], enzyme kinematics [holmberg1982practical] or biochemical reactors [strogatz2024nonlinear], and references therein. In general, identification of such systems can be challenging [evans2002identifiability]. In addition to this, problems in control can also give rise to system of differential equations in which the function $f$ can have discontinuities like finite jumps (switching on or off at certain time) or sudden changes in the behavior of the system [chappell1990global]. Once during a test or simulation if a point of discontinuity is located, then our goal is to design the learning algorithm in such a way that it takes care of such discontinuous points by associating the symbol $f$ with suitable Hilbert space (See Figure 1). It is our desire to explore this avenue and start our quest of deep analysis towards this direction. Although, it is always not possible to figure out all such points during simulations, nonetheless we will start our investigation assuming that such points can be traced. The following items are the main contribution of this paper;
@@ 3.1 Reproducing Kernel Hilbert Space: RKHS
It is the closed subspace of $L^{2}(\mathbb{T})$. Vice versa, we can define the functions in the Hardy-Hilbert space $H^{2}(\mathbb{D})$, as the integral of the functions in $H^{2}(\mathbb{T})$ against the Poisson Kernel [fricain2016theory]. The reproducing kernel of $H^{2}(\mathbb{D})$ is known as the Szego Kernel and it is given by
@@
Kernel functions interact nicely with many classical function theoretic operators. For example, if $f:X\to\mathbb{C}$ is the symbol for a densely defined multiplication operator, $M_{f}:\mathcal{D}(M_{f})\to H$, with $\mathcal{D}(M_{f}):=\{g\in H:fg\in H\}$ and $M_{f}g:=gf$, then $M_{f}^{*}k_{x}=\overline{f(x)}k_{x}$. That is kernel functions are eigenfunctions for the adjoints of densely defined multiplication operators [paulsen2016introduction]. Similarly, the adjoint of a densely defined Koopman or composition operator, $\mathcal{K}_{\phi}$, on a kernel function is described as $\mathcal{K}_{\phi}^{*}k_{x}=k_{\phi(x)}$.
@@ 3.2 Liouville Operator and Occupation Kernel
It is easy to check that $A_{f}$ is a linear operator. Being a differential operator, $A_{f}$ is expected to be unbounded operator over $H^{2}(\mathbb{D})$, but as differential operators over RKHS’s consisting of continuously differentiable functions, it is always closed [rosenfeld2024occupation], hence $A_{f}$ is a closed operator over $H^{2}(\mathbb{D)}$.
@@
The above relation can be used to find finite approximation of the Liouville operator. For more information, we refer to the papers [rosenfeld2024occupation].
@@ 4 Restricted Liouville Operator
In this paper, we will pose the above problem inside $H^{2}(\mathbb{D})$, by taking $f$ as a symbol which is not analytic or smooth in nature. The following theorem from [russo2022liouville], explains the behavior of the symbol $f$ of the operator $A_{f}$, when $\overline{D(A_{f})}=H^{2}(\mathbb{D})$.
@@
Then it can be easily proved that $B_{j}\in H^{\infty}(\mathbb{D})$ and $|B_{j}(e^{i\theta})|=1$ for all $e^{i\theta}\in\mathbb{T}$. In literature, the function $B_{j}$’s are known as the finite Blaschke factors. One can easily show that $\|B_{j}g\|_{H^{2}}=\|g\|_{H^{2}}$ for all $g\in H^{2}$, and hence multiplication operator $T_{B_{j}}:H^{2}\rightarrow H^{2}$ is an isometry, which implies that the range space $T_{B_{i}}H^{2}=B_{i}H^{2}$ is an closed subspace of $H^{2}$ (see, [lata2022multivariable]).
@@
Moreover, it is easy to see that $B\in H^{\infty}(\mathbb{D})$ and $|B(e^{i\theta})|=1$ for all $e^{i\theta}\in\mathbb{T}$ (see, [fricain2016theory]). Now, we are in the position to construct an Hilbert space which captures the singularities of the function $f$. Define the set
@@
Then $B^{2}H^{2}$ is a closed subspace of $H^{2}$ with $\{B^{2}z^{n}\}_{n=0}^{\infty}$ as its orthonormal basis [lata2022multivariable]. Moreover, according to [garcia2018finite], the space $B^{2}H^{2}$ is an RKHS with the reproducing kernel defined as
@@ 6 Spectrum of Restricted Liouville Operator
Note, the symbol $f(z)=1/z$, has no zeroes inside the disk and has a pole inside the disk. On the other hand, according to Proposition 5 in [russo2022liouville], if $f$ is an analytic function with no zeroes in a neighborhood of the closed disk, then every $\lambda\in\mathbb{C}$ is in the point spectrum with $g(z)=C~\text{exp}~\Big(\int_{0}^{z}\frac{\lambda}{f(w)}dw\Big)$ as the corresponding eigenfunction. We can further improve Proposition 6.1, in the following way.
@@
    Spectrum of $A_{1/z}$ is continuous. Combining Proposition 6.1 and Theorem 6.2, we can conclude that the spectrum of the operator $A_{1/z}$ is whole complex plane, i.e., $\sigma(A_{1/z})=\mathbb{C}$. This is in contrast with the case $A_{z}$ where $\sigma(A_{z})=\mathbb{C}$, where the spectrum consists of only eigenvalues [Proposition 5, [russo2022liouville]]. Hence, the spectrum of the operator $A_{f_{|B^{2}H^{2}}}$ needs further investigations.